I asked this earlier. Thanks Ed for replying. I could npt get the site to open where I could say thanks. I got the same answer as you did on the first problem I submitted, however like I stated before it was not an option. Here are all the options I was given. PLEASE help me!!
-------------------------------------------------------------------------
Add. Express your answer in simplest form. (8/s)+(8/(s^2))

A)16/s^2 B)8s+8/s^2 C)8s+8/s D)16/s
-------------------------------------------------------------------------
During rush hour, Bill can drive 15 miles using the side roads in the same time that it takes to travel 10 miles on the freeway. If Bill's rate on the side roads is 8 mi/h faster than his rate on the freeway, find his rate on the side roads.
I answered....rate on side roads = 24, correct?

Answer by edjones(7569)

(8s+8)/s^2 is ok.
.
d=rt distance = rate*time
f=freeway rate
15=(f+8)t...t=15/(f+8)
10=ft...t=10/f
10/f=15/(f+8)
10(f+8)=15f
10f+80=15f
5f=80
f=16
16+8=24 on side roads
UR RIGHT!!!
.
Ed